Sometimes both equations will need to be multiplied to obtain opposite coefficients
EXAMPLE
3X + 5Y =6
-4X + 2Y =5

Found 2 solutions by checkley75, stanbon:
Answer by checkley75(3666)

MULTIPLY THE FIRST EQUATION BY 4 WE GET
12X+20Y=24 NOW MULTIPLY TH ESECOND EQUATION BY 3 WE GET:
-12X+6Y=15 NOW WE ADD TOP GET RID OF THE X TERMS THUS:
--------------------
26Y=39
Y=39/26
Y=3/2 ANSWER.
NOW SUBSTITUTE 3/2 FOR Y & SOLVE FOR X IN EITHER EQUATION.
3X+5(3/2)=6
3X15/2=6
3X=6-15/2
3X=(12-15)/2
3X=-3/2
X=-3/2*1/3
X=-1/2 ANSWER.

Answer by stanbon(75887)

3X + 5Y =6
-4X + 2Y =5
-------------
Let's say you want to get rid of the X-terms:
Multiply 1st equation by 4 so you get a 12X term:
Multiply 2nd equation by 3 so you get a -12x term:
----------
12x + 20y = 24
-12x+ 6y = 15
-----------
Now, add the two equations to get:
26y=39
y=39/26
-----------
Substitute that value of y into one of the equations so you can solve for x:
3x +5(39/26) = 6
x+5(13/26) = 2
x+5/2 = 2
x= -1/2
===========
Cheers,
Stan H.